# A Survey on Federated Analytics: Taxonomy, Enabling Techniques, Applications and Open Issues

### v1

@@ VIII Database operations
In Table VIII, we summarize our surveyed federated query solutions, regarding their supported query types, privacy models, cryptography techniques, and maximal supported clients.
TABLE VIII: Summery of federated query solutions.
Reference Query type Privacy Cryptography technique Maximal client
SMCQL [46] General Encryption Garbled circuit 2
Conclave [47] General Encryption Secret sharing 3
Senate [50] General Encryption (adversary model) Garbled circuit 16
Saqe [49] General Encryption & DP Secret sharing 2
Orchard [53] General Encryption & DP Homomorphic encryption $>$1000
Arboretum [55] General Encryption & DP Homomorphic encryption $>$1000
Secure Yannakakis [48] Free-connex join-aggregate query Encryption Private set intersection (based on garbled circuit), secret sharing, & oblivious extended permutation 2
Zhang et al. [51] Skyline query Encryption Private set intersection (based on homomorphic encryption and garbled bloom filter) 10
Hu-fu [52] Spatial query Encryption Secret sharing 10
Mycelium [54] Graph query Encryption & DP Homomorphic encryption & onion routing $>$1000
IX Model-based applications
The model-based formulations and algorithms have gained great success in handling complex data analytics tasks. FA-based methods become an effective solution for model-based applications when the privacy-sensitive data stored in distributed clients are required for training the model. In this section, we review several well-developed model-based applications and their existing FA solutions, including linear regression model (Section IX-A), Bayesian model (Section IX-B), matrix factorization model (Section IX-C), multi-armed bandits model (Section IX-D), and unsupervised representation learning model (Section IX-E).
IX-A Linear regression
Linear regression is a classical model-based data analytics problem. It discovers the linear relationship between features and observations. Consider a dataset $\mathcal{N}$ with $n$ data points, for a data point with index $i$, $X_{i}$ is its vector in the form of a $k$-length vector, and $y_{i}$ is its observation in the form of a scaler. Denote $\boldsymbol{X}=[X_{1},...,X_{n}]$, and $\boldsymbol{y}=[y_{1},...,y_{n}]$, the model of linear regression assumes the linear relationship between $X_{i}$ and $y$, i.e.,
$\boldsymbol{y}=\boldsymbol{\beta X}+\boldsymbol{\epsilon},$ (10)
where $\boldsymbol{\beta}$ is the parameters of the linear regression model in the form of a $k$-length vector, and $\boldsymbol{\epsilon}$ denotes the independent noise for each data point. The objective of the linear regression problem is to derive a set of parameters $\boldsymbol{\beta}$ that can fit the relationship well, e.g., that minimize the square error (MSE), i.e.,
$\min_{\boldsymbol{\beta}}\sum_{i=1}^{n}(\boldsymbol{\beta X}_{i}-y_{i})^{2},$ (11)
or minimize the regularized square error, well known as ridge regression, i.e.,
$\min_{\boldsymbol{\beta}}\sum_{i=1}^{n}(\boldsymbol{\beta X}_{i}-y_{i})^{2}+||\boldsymbol{\beta}||^{2}$ (12)
In the federated setting, the whole dataset $\mathcal{N}=\{\boldsymbol{X},\boldsymbol{y}\}$ is distributed among multiple clients. For a better explanation, we construct $\mathcal{N}$ to be a matrix with $n$ rows and $k+1$ columns, and the $i$-th column is constructed by the $k$ entries of $X_{i}$, followed by $y_{i}$, so that each row of the matrix $\mathcal{N}$ exactly represents all information of one data point. Federated linear regression can be categorized into horizontal federated linear regression and vertical federated linear regression, based on how the data $\mathcal{N}$ is distributed among clients. In the horizontal setting, each client possesses a subset of rows of $\mathcal{N}$, i.e., each client possesses the complete information of a part of the data points. In the vertical setting, each client possesses a subset of columns of $\mathcal{N}$, i.e., each client possesses incomplete information (some dimensions) of all the data points. The two settings of federated linear regression introduce different challenges, resulting in varying techniques utilized.
In [127], the authors propose a horizontal federated linear regression solution, termed federated data analytics (FDA). It utilizes the fact that in horizontal federated linear regression, as a client possesses the complete information of some data points, it can independently evaluate the error of these data points, and derive the gradient of the linear regression parameters $\boldsymbol{\beta}$. FDA is based on a hierarchical architecture, i.e., the client’s local sets of linear regression parameters (low hierarchy) and a set of shared parameters held by the server (high hierarchy) are all updated based on the information learned from each other. The authors then provide adaptations of the linear regression model on various downstream tasks, including hypothesis testing, uncertainty quantification, and variable selection. In [128], another horizontal federated linear regression solution LDP-IHT is proposed, also applying the idea of gradient transmission. It applies the LDP scheme on the uploaded gradient for better privacy preservation. The authors then investigate the harm to data utility (represented by error bound) of applying LDP on the high-dimensional gradients and propose the sequential interactive algorithm LDP-IHT to reduce the harm. In [129], another solution is proposed, where the computation of horizontal federated linear regression is encrypted by homomorphic encryption and garbled circuit. This work takes a stricter setting, that one client only possesses features and observation of a single data point. DP-CFL [130] is a horizontal federated linear regression solution that transmits the coded version of the local features and observations to the server for privacy preservation. The coding procedure is designed to multiply the data matrix with a random matrix sampled from the Gaussian distribution so that DP guarantee is enforced by DP-CFL. The aggregated coded data is the intermediate result to compute gradient on $\boldsymbol{\beta}$, so the server can learn the optimized $\boldsymbol{\beta}$ based on the aggregated coded data.
In the vertical federated linear regression setting, an extra challenge is introduced, as a client can no longer evaluate any single data point, and clients need to collaborate to derive the full information of any data point with privacy preservation. In [131], an early vertical federated linear regression solution is proposed, where Powell’s algorithm is applied to let clients optimize their parts of $\boldsymbol{\beta}$, and a secure summation protocol encrypts the intermediate computation results. In [132], the authors propose an improved solution for vertical federated linear regression. They design a novel MPC protocol combining the garbled circuit and a tailored protocol for secure computation of the inner product. They propose a conjugate gradient descent-based solver for linear regression problems, which is more suitable for MPC schemes. Another vertical federated linear regression solution, proposed in [133], gains a major advantage: it only utilizes linear homomorphic encryption to achieve security, with the absence of other computation-intensive cryptography tools. In [134], the authors investigate the heavy communication overhead of the feature and observation sharing among clients and introduce the idea of performing computation on coreset, a small subset of data points that can derive an accurate enough linear regression result, instead of the whole dataset. The authors then propose an important sampling method to derive the coreset. The coreset solution is also applied in the federated clustering problem, which is introduced in the corresponding part of this survey.
In [135], the power of edge computing is leveraged to accelerate the vertical federated linear regression computation, where an erasure code-based scheme is utilized to allocate the client workload to the edge workers, and homomorphic encryption is utilized for privacy guarantee.
IX-B Bayesian model
Bayesian analysis is a powerful probabilistic tool to handle the prior and posterior distributions of complex probabilistic events. It has many important downstream topics. In Bayesian optimization, the characteristics of an unknown function are evaluated based on a set of observations (data points), while the function itself is considered a black box during the procedure. In Bayesian network learning, the casual relationship among a number of random variables is evaluated based on the observations. The FA version of Bayesian analysis is naturally motivated as the scenario is considered that the observations, necessary for any Bayesian analysis task, are held by different clients. The FA-based Bayesian analysis problem targets collaboratively deriving the result of Bayesian analysis with the observations held by the clients while preserving the privacy of the observations.
In [136], the authors propose a federated Bayesian optimization solution. It designs an FA version of the Thompson sampling to resolve the federated Bayesian optimization problem and proposes the random Fourier feature that enables the sharing of useful information for Bayesian optimization without transmitting raw data. In [137], the previous federated Bayesian optimization model is improved, where the privacy preservation is reinforced by DP, and a novel distributed exploration technique accelerates the exploration phase of the algorithm. In [19], a digital twin-assisted FA scheme is proposed for Bayesian optimization on distribution estimation. It utilizes an FA version of MCMC combined with a rejection–acceptance sampling technique and a delayed rejection technique. In [85], the observation values of the clients in the federated Bayesian optimization system are protected with LDP satisfaction based on the Laplace mechanism. In [138], a trend prediction problem is formulated by the Bayesian optimization problem and is resolved in an FA model. In that model, observations in each client are summarized by a distribution described by a feature vector. Trend prediction is conducted by analyzing the distributions over a time series. To aggregate the distributions, secure aggregation is performed on the feature vectors based on secret sharing. In [139], a solution for federated Bayesian network structure learning is proposed. The problem is formulated as continuous optimization and is resolved by the FA version of the ADMM algorithm so that only model parameters are exchanged.
IX-C Matrix factorization
TABLE IX: Summery of federated matrix factorization solutions.
Reference Type Separation Privacy Extra features
Dolui et al. [140] Embedding Horizontal DP & Data anonymization Application on a recommender system
FedMF [56] Embedding Horizontal Homomorphic encryption Gradient-based mathod
F2MF [141] Embedding Horizontal DP Fairness concern
Li et al. [57] Embedding Horizontal & Multiple rows & Discrete entries DP & Secure aggregation & Gradient clipping Abstracted computation paradigm for three data separations
Gao et al. [142] Embedding Horizontal & Discrete entries Null Analysis on privacy attack models; Auxiliary data setting
Grammenos et al. [143] SVD Multiple rows (streaming data) DP Time-independent & asynchronous
FedSVD [58] SVD Multiple rows Secure aggregation Lossless SVD results
FedPower[144] SVD Horizontal DP Orthogonal Procrustes transformation; Power method
Hartebrodt et al. [145] SVD Multiple rows Null No eigenvector uploading; Preventing information leakage via smart initialization and data approximation
(a) Horizontal
(b) Multiple rows
(c) Discrete entries
Figure 8: Illustration of different data separation settings in federated matrix factorization
Matrix decomposition, or matrix factorization, refers to the procedure of decomposing one data matrix $M$ into multiple output matrics so that the product of the output matrics is exactly (or approximates) the data matrix. The matrix decomposition problem has many variations, categorized by the required characteristics of the output matrics.
Matrix embedding factorization [^4] Many recommender system researchers directly use the term “matrix factorization” to denote matrix embedding factorization considers a data matrix $M$ with shape $m\times n$ ($m$ rows and $n$ columns), and derives two output matrices $A$ with shape $m\times k$ and $B$ with shape $k\times n$, i.e.,
$M[m\times n]=A[m\times k]B[k\times n].$ (13)
Matrix embedding factorization has many important applications in recommender systems and collaborative filtering, as it directly models the recommendation problems of production environments. Consider a recommender system setting with $m$ consumers (users) and $n$ products for advertisement (items), and the preference of user $i$ towards item $j$ is recorded by the $(i,j)$ entry of the data matrix $M$. The matrix embedding factorization algorithms output the user embedding matrix $A$, and the item embedding matrix $B$, where each row of $A$ is the $k$-length embedding of one user, and each column of $B$ is the $k$-length embedding of one item. With matrix embedding factorization, each user and item can be summarized by a fixed-length vector, representing their characteristics in the user-item interaction. As $k$ is usually set to be much smaller than $m$ and $n$, the equation $M=AB$ can not be guaranteed to have a solution. Therefore, the algorithm designers instead output $A$ and $B$ that minimize the square error in deriving user-item preferences, i.e.,
$\min_{A,B}\sum_{M(i,j)\in M}(M(i,j)-A(i,:)B(:,j))^{2},$ (14)
or minimize the normalized square error, i.e.,
$\min_{A,B}\sum_{M(i,j)\in M}(M(i,j)-A(i,:)B(:,j))^{2}+\lambda(||A||^{2}+||B||^{2}).$ (15)
The aforementioned practical scenario setting of matrix embedding factorization naturally yields the need to solve it in a federated way. Existing federated matrix factorization solutions can be roughly categorized into three categories, based on how the data matrix $M$ is separated among clients: horizontal setting, multiple rows setting, and discrete settings. In the horizontal setting, each client exactly possesses one row of the $M$. In the multiple rows setting, each client possesses multiple rows of $M$ [^5] The term “row” in the two settings can be replaced by “column”, without loss in correctness. In the discrete entries setting, one client cannot obtain a complete row. Information of any complete row is separated by multiple clients.
Horizontal federated matrix factorization originates from the setting that the FA system treats the users as the clients which exactly hold one row of the data matrix $M$ [140, 56, 141]. [^6] This is equivalent to the case that each client holds one column of $M$. In [140], the authors propose a federated matrix factorization algorithm that considers the common case that each client exactly holds all ratings of one user (one row of $M$). Their proposed algorithm lets clients compute the gradients on $A$ and $B$, and transmit the gradients instead of raw data. In [56], another federated matrix factorization solution FedMF is proposed considering a similar problem setting. It also applies gradient transmission, and utilizes homomorphic encryption to prevent the privacy risk caused by exposing the gradient to the server. In [141], the user group fairness – fairness of recommendation accuracy among different user groups – is considered in the federated matrix factorization problem. The authors design a DP scheme to communicate the matrics needed for computing the group fairness.
Other data separation settings of federated matrix factorization are also investigated. In [57], in addition to the aforementioned horizontal setting, the authors consider the setting where each client possesses ratings of a subset of users (multiple rows setting), and the setting where each client possesses all users’ rating on a subset of items (discrete entries setting). The authors then propose FMF as a unified framework to resolve all these data separation settings, with the embedding clipping technique to ensure DP. In [142], the authors consider the setting where each client possesses partial item ratings on a subset of users (discrete entries setting), and the setting where some clients possess auxiliary data of user-item interaction. The authors then discuss the privacy threats and available privacy-preserving approaches in these settings.
Singular value decomposition (SVD) is another important variation of matrix factorization. For a data matrix $M$ with shape $m\times n$, SVD decomposes $M$ into the product of three output matrics $U,V$ and $\Sigma$, so that
$M=U\Sigma V,$ (16)
where $U$ with shape $m\times m$ is a complex unitary matrix, $\Sigma$ with shape $m\times n$ is a rectangular diagonal matrix with non-negative values (singular values of $M$) on the diagonal, and $V$ with shape $n\times n$ matrix with its conjugate transpose $V^{*}$ to be a complex unitary matrix.
As a fundamental matrix factorization technique, SVD has many important applications in data analytics, such as recommender systems. Here we make a similar example to that of matrix embedding decomposition. Consider that $M$ stores user-item rating data with $m$ users and $n$ items. The SVD results are also able to derive vertex-type representations of the users and items. Essentially, each row of $U$ is the embedding of one user, and each row of $V^{*}$ is the embedding of one item.
Federated SVD solutions are then proposed to tackle the scenarios when the data in $M$ is held by distributed clients. These solutions output the SVD results while preserving the privacy of client-held input data. In [143], the authors propose a federated principle component analysis algorithm, whose outputs are exactly approximations of the SVD results. The authors consider the scenario where each client possesses some rows of the data matrix $M$. The local data of each client is generated in a streaming fashion, which leads to the challenge of limited local storage. The proposed algorithm includes tailor-designed local computation and aggregation schemes and satisfies DP as a formal privacy guarantee. In [58], FedSVD is proposed as another federated SVD solution. It relies on a trusted authority that distributes masks to the clients. The clients add masks to their data and upload the masked data to the server with secure aggregation. The masks are delicately designed for SVD so that users can easily unmask the computation results received from the server. With FedSVD, SVD results can be derived accurately without privacy preservation.
In [144], FedPower is proposed as a federated eigenspace estimation (part of SVD procedure) algorithm. It designs a distributed version of the power method, a classical eigenspace estimation scheme, and leverages the orthogonal Procrustes transformation to improve the aggregation phase of the federated power method, to resolve the orthogonal ambiguity. Gaussian noise is added to the possessed matrix in each iteration to enforce Rényi DP. In [145], a federated SVD solution is proposed for genome-wide association studies. Their solution is based on the Gram-Schmidt algorithm. The authors emphasize the advantages of genome-wide association studies, as the computation overhead is independent of the sample size, only depending on the size of features.
(a) Server-client architecture with global model
(b) Server-client architecture with personalized model
(c) Fully decentralized architecture
Figure 9: Illustration of three types of federated MAB settings
TABLE X: Summery of federated MAB solutions.
Reference Architecture Privacy Extra features
FMAB [59] Server-clinet Null Global model; Reward heterogeneity
PF-MAB [60] Server-clinet Null Personalized model; Reward heterogeneity
Li et al. [61] Both DP Global model; Privacy-regret trade-off
Zhu et al. [62] Decentralized DP Personalized model
FedUCB [63] Both Federated DP Global model; Contextual bandit
IX-D Multi-armed bandits
The multi-armed bandits (MAB) model considers a system with $k$ arms, each with a fixed distribution of rewards, and an agent successively pulling the arms. Each time the agent selects an arm and pulls it, a stochastic reward determined by its corresponding reward distribution is returned to the agent. The agent aims at maximizing the reward it received, or formally speaking, minimizing the cumulative regret $\rho$:
$\rho=T\mu^{*}-\sum_{t=1}^{T}r_{t}.$ (17)
The aforementioned definition $\rho$ indicates the cumulative regret after $T$ pulls, where $\mu^{*}$ indicates the maximal reward mean among all arms, and $r_{t}$ indicates the received reward in the $t$-th pull. An MAB algorithm performs decision-making in selecting the arm to be pulled in each round. The algorithm handles the famous exploration-exploitation tradeoff: whether the agent should select an underexplored arm with the potential to have a high reward mean, or should exploit an arm that has a relatively high reward mean based on previous trials.
As the effectiveness of an MAB algorithm highly relies on gathering and analyzing previous trial results, the federated MAB model emerges to enable multiple clients (MAB agents) to collaboratively improve their own MAB performances. The federated MAB model assumes that multiple clients are having their MAB models with a shared set of arms. By sharing the trial results, the MAB training can be greatly improved.
There exist at least three kinds of federated MAB setting, as demonstrated in Fig. 9: 1) Server-client architecture with global model, where a single server model is optimized by aggregating the trial results from all clients; 2) Server-client architecture with personalized model, where each client processes its own MAB model, and optimizes its local model by leveraging both trial results from itself and trail results of other clients received from the server; 3) Decentralized architecutre, where personalized models are also trained by the clients, but no centralized server exists to aggregate all trial results. Instead, each client can only receive trial results from a portion of clients via client-to-client communication.
In [59], a federated MAB solution is proposed. It considers the scenarios that the server would like to learn the global bandit model, while the MAB model of the clients is diverged from the global model regarding the arm reward means. To tackle the challenge of diverged MAB models, in addition to the basic federation scheme which transmits the trial results between server and clients, the authors propose a novel confidence bound-based scheme, named Fed2-UCB, to handle the uncertainty of sampling clients with the heterogeneous arm reward means. In [60], another federated MAB solution is proposed to handle a different federated MAB setting. The new solution also handles the problem of heterogeneous reward means. In addition to the global bandit model, in [60], each client would like to train its personalized MAB model that fits its local environment. By carefully leveraging the knowledge from other clients, their proposed federated MAB algorithm, PF-MAB, achieves good regret bounds for each client, despite that the information from other clients might be misleading. In [61], the authors consider enhancing privacy preservation in transmitting the trial information. They apply DP to the uploads and modify the UCB algorithm of the MAB model to handle the uncertainty introduced by DP. Their system is deployed in both the server-client architecture and the fully decentralized architecture, where clients independently communicate with other clients and share information about the trial results. In [62], the authors further investigate the fully decentralized federated MAB models. The proposed Gossip-UCB algorithm tackles the coupling effect of client gossiping, considers the effect of biased client rewards, enforces DP, and provides a theoretical guarantee on regrets. In [63], the authors investigate the federation of contextual bandits, a variation of MAB with parameterized trials and arms. In their proposed scheme, the Gram matrix of actions and reward vector, which are the information about the regression of rewards, are transmitted. The proposed FedUCB algorithm, based on the LinUCB algorithm for centralized contextual bandits, handles both the server clients and decentralized architectures and enforces DP for better privacy preservation.
IX-E Unsupervised representation learning
Unsupervised representation learning refers to the procedure of deriving expressive representations (features) for data in various structures, where the data are usually unlabelled. It converts the raw data with complex structure and high dimensionality into vectors with limited dimensions, which are feasible for many downstream tasks. Unsupervised representation learning is also applied in federated scenarios, where the data stored in multiple clients are given their features in a common scope. The federated representation learning problem is often resolved in the FA paradigm, where clients upload insights about their local data and features, and the server aggregates the insights to derive a comprehensive view of the data distribution and feature space. Many representation learning algorithms utilize deep learning, and their federated versions utilize FL. However, FA is still needed in these algorithms as the complex information exchange cannot be accomplished by simply averaging the parameters of neural networks.
In [146], a federated representation learning algorithm that utilizes both FA and FL is proposed. It realizes unsupervised representation learning via contrastive learning. It trains a feature learning model via FL and utilizes FA to exchange local features with other clients. With the knowledge of global feature distribution, contrastive learning can be properly performed to derive distinguishable features, even if the data within each client is highly skewed. In [147], an unsupervised federated representation learning method FedCA is proposed. In addition to the contrastive learning model trained by FL, FedCA proposes a dictionary module trained by FA to align the learned representations among clients with non-IID data. The dictionary model is first derived locally by adding the normalized representations of local data samples and then aggregated by the server to derive the global dictionary. FedWalk [148] is an FA-based solution for graph embedding analytics. It derives expressive representations for the vertexes in a large graph, which has a complex and uncertain data structure. FedWalk considers the vertex-separated federated setting, where each client represents one vertex in the graph, and only possesses neighborhood information of the vertex. It performs the FA version of a random walk, a classical graph embedding learning algorithm.
X Assisting federated learning
Figure 10: A typical workflow of FL assisted by FA.
@@ X Assisting federated learning
We categorize the solutions of FA-assisted FL based on when FA takes effect, as illustrated in Fig. 10. We summarize the whole procedure into three phases: the preprocessing phase, where the clients are registered and evaluated by the server; the training phase, where the FL model is transmitted between server and clients, trained by the clients locally, and aggregated by the server, for multiple rounds; the postprocessing phase, where the performance of the derived FL model is then evaluated. Strictly speaking, FL only includes the training phase, and the assisting FA schemes can be executed in the preprocessing phase, in the postprocessing phase, and sometimes in the training phase parallel with FL. Fig. 10 illustrates the involvement of FL and FA in the whole procedure with colors.
@@
In the following parts, we will introduce some classes of FA schemes to assist FL. In Section X-A, we investigate FA schemes that help FL to select proper participating clients, which executes on the preprocessing phase and client selection step of the training phase. In Section X-B, we investigate FA schemes that help FL model personalization via conducting clustering on the clients, which executes on the preprocessing phase and model aggregation step of the training phase. In Section X-C, we investigate FA schemes to evaluate the performance of the FL model, which executes in the postprocessing phase.
X-A Preprocessing: Client selection in FL
In FL, only part of the clients will participate in each round of FL, due to the limits of device availability and system capacity. Many pieces of research show that clients provide different benefits/contributions to the model training determined by the data size and quality of the clients. Therefore, wisely selecting the participating clients in each round then has the potential to obtain a better FL performance, compared to the default setting of random selection. Selecting the proper participating clients requires analyzing the characteristics of the client data, which naturally requests the power of FA.

### v3

@@ VIII Database Operations
In Table IX, we summarize our surveyed federated query solutions, regarding their supported query types, privacy models, cryptography techniques, and maximal supported clients.
Lessons learned. Federated query solutions are crucial for enabling privacy-preserving data analytics across distributed clients, but they face significant challenges in computation overhead due to the use of cryptographic tools like MPC. Optimizing computation time is essential, with researchers exploring various methods such as lightweight cryptographic techniques, query graph optimization, and approximations. While general solutions support a wide range of queries, they often struggle with scalability, whereas specific query solutions can achieve better performance by focusing on particular types of queries. Scalability and performance improvements highlight the importance of adapting federated query systems to the specific needs of large-scale environments and specialized queries.
IX Assisting Federated Learning
TABLE X: Summary of assisting federated learning.
Reference Data analytics task Note
FedACS [14] Preprocessing: Client quality evaluation for FL Selecting clients with lower data heterogeneity based on their gradients
FAVOR [92] Preprocessing: Client quality evaluation for FL Using a reinforcement learning model to select clients based on their gradients to maximize validation accuracy
Oort [87] Preprocessing: Client quality evaluation for FL Selecting clients with higher training losses
Cho et al. [84] Preprocessing: Client quality evaluation for FL Prioritizing high-loss clients for participation
CMFL [85] Preprocessing: Client quality evaluation for FL Preventing gradient uploads of low-quality clients
Sattler et al. [15] Preprocessing: Personalize FL models via client clustering Grouping clients with similar data distributions
Briggs et al. [89] Preprocessing: Personalize FL models via client clustering Agglomeratively clustering clients after training a global model
Liang et al. [88] Preprocessing: Personalize FL models via client clustering Decomposing and consolidating extreme clusters
IFCA [90] Preprocessing: Personalize FL models via client clustering Assigning clients to clusters where they achieve the lowest local loss on distributed models
COMET [91] Preprocessing: Personalize FL models via client clustering Incorporating knowledge sharing with a regularization term during local model training
FedSoft [86] Preprocessing: Personalize FL models via client clustering Using proximal loss to align client training with multiple cluster models
Oort [87] Postprocessing: Model evaluation Automatically adjusting the number of participating clients using confidence bounds derived from Hoeffding’s inequality
Figure 10: A typical workflow of FL assisted by FA. FA is applied in the postprocessing phase of the FL pipeline. The preprocessing phase, as well as the client selection and model aggregation steps in the training phase, includes the involvement of both FL and FA.
In this section, we focus on the fourth class of data analytics applications following our taxonomy: assisting federated learning. We consider three scenarios where FL can gain benefit from FA-empowered analytics, and introduce them separately in the following part.
@@ IX Assisting Federated Learning
We categorize the solutions of FA-assisted FL based on when FA takes effect, as illustrated in Fig. 10. We summarize the whole procedure into three phases: the preprocessing phase, where the clients are registered and evaluated by the server; the training phase, where the FL model is transmitted between server and clients, trained by the clients locally, and aggregated by the server, for multiple rounds; the postprocessing phase, where the performance of the derived FL model is then evaluated. Following the definition of the majority of FL research, FL usually includes the training phase and a part of the preprocessing phase, and the assisting FA schemes can be executed in the preprocessing phase, in the postprocessing phase, and sometimes in the training phase parallel with FL. Fig. 10 illustrates the involvement of FL and FA in the whole procedure with colors.
The FA schemes for FL assistance usually do not apply any privatization methodology other than “hiding raw data”. Specifically, since many FA schemes recycle the client uploads in the host FL mechanisms as FA insights, they essentially inherit the privacy preservation of the FL schemes. For example, when the host FL schemes utilize LDP in the uploaded gradients, the assistance FA schemes also gain an LDP guarantee if they utilize the FL gradient as insights (it is common in FA-assisted FL).
@@
In the following parts, we will introduce some classes of FA schemes to assist FL. In Section IX-A, we investigate FA schemes that help FL to evaluate the data quality of clients, and further select proper participating clients, which executes on the preprocessing phase and client selection step of the training phase. In Section IX-B, we investigate FA schemes that help FL model personalization via conducting clustering on the clients, which executes on the preprocessing phase and model aggregation step of the training phase. In Section IX-C, we investigate FA schemes to evaluate the performance of the FL model, which executes in the postprocessing phase. Table X summarizes the FA schemes covered by this survey in assisting FL applications.
IX-A Preprocessing: Client quality evaluation for FL
Being consistent with conventional model training, the quality of training data is a critical factor in influencing the outcome model performance in FL. When it comes to the scenario of federated data, the quality of data in different clients varies, providing different benefits (or even harms) to the FL model. The different quality of client data may originate from the data intrinsically, like different correctness of labels, and different scales of noise, or from the overall FL system, like different heterogeneity of data distribution compared to the global one (aka. non-IID issue). Therefore, evaluating the quality of client data in the preprocessing phase becomes a profitable procedure. The results of client quality evaluation can be utilized for many FL system optimizations. In some FL research, the client quality evaluation is employed to improve the client selection of FL, rather than the random selection baseline. In other words, the uploaded gradients from different clients are processed by the server before aggregation, through reweighting or trimming, based on the client evaluation results.